# Regularized Q-Learning with Linear Function Approximation

### v1

@@ 2.2 Linear Function Approximation
In environments with large state spaces, such as continuous spaces, maintaining a lookup table to store and process values for all possible state-action pairs is impractical. Function approximation techniques are prevalent as they offer an effective solution to this problem by allowing agents to approximate the values using significantly fewer parameters. In this study, we focus on leveraging linear function approximation which approximates the value functions linearly in the given basis vectors. The basis vectors $\phi_{i}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$, $i=1,2,\dots,d$ are, without loss of generality, linearly independent. We denote $\phi(s,a)=\left[\phi_{1}(s,a),\phi_{2}(s,a),\dots,\phi_{d}(s,a)\right]^{\top}\in\mathbb{R}^{d}$ for all $(s,a)$ and let $\Phi\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times d}$ defined by $\Phi=\left[\phi(s_{1},a_{1}),\dots,\phi(s_{|\mathcal{S}|},a_{|\mathcal{A}|})\right]^{\top}$. We approximate state-action function as $\hat{Q}_{\theta}\triangleq\Phi\theta$, where $\theta\in\mathbb{R}^{d}$ is the parameter vector. We also use $\hat{Q}_{\theta}(s,\cdot)=\phi(s,\cdot)^{\top}\theta$ to denote the vector of the approximated state-action values of all actions given the state $s\in\mathcal{S}$, where $\phi(s,\cdot)=\left[\phi(s,a_{1}),\dots,\phi(s,a_{|\mathcal{A}|})\right]$. Suppose $\mu$ is a distribution that exhibits full support across the state-action space, let $D_{\mu}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}||\mathcal{A}|}$ be a diagonal matrix with the distribution $\mu$ on its diagonal. Furthermore, we define the weighted $\ell_{2}$-norm with respect to the distribution $\mu$ as $\|x\|_{D_{\mu}}\triangleq\sqrt{x^{\top}D_{\mu}x}$. Let $\Pi_{D_{\mu}}$ be the projection operator which maps a vector onto the span of the basis functions with respect to the norm $\|\cdot\|_{D_{\mu}}$; it is given by $\Pi_{D_{\mu}}=\Phi\left(\Phi^{\top}D_{\mu}\Phi\right)^{-1}\Phi^{\top}D_{\mu}$ under the assumption that $\Phi^{\top}D_{\mu}\Phi$ is a positive definite matrix.
@@ 3.1 Smooth Truncation Operator
In addition, we let $\mathcal{B}_{\tau,\delta}$ be the smooth truncated optimal regularized Bellman operator which is given as, for an arbitrary mapping $X:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$:
$\mathcal{B}_{\tau,\delta}X(s,a)\triangleq R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\left[\hat{G}_{\tau,\delta}\left(X\left(s^{\prime},\cdot\right)\right)\right],$
@@
3.2 Bi-level Formulation
Before presenting the problem formulation, we outline some assumptions that are crucial for our analysis. Let $\pi_{\mathrm{bhv}}$ be the behavioral policy used to collect data in regularized MDP $\mathcal{M}_{\tau}$ and we adopt the following assumption.
Assumption 3.1
Assumption 3.1 guarantees the ergodicity of the Markov chain, which in turn ensures that it converges to its stationary distribution at a geometric rate. Furthermore, Let $\mathcal{D}$ be the stationary distribution of the transition tuples induced by $\pi_{\mathrm{bhv}}$. Consequently, $\mathcal{D}=\mu_{\mathrm{bhv}}\otimes P$, where $\otimes$ denotes the tensor product between two distributions. We use the term $\mathcal{D}$ to also refer to the marginal distribution of state-action pairs, i.e., $\mathcal{D}(s,a)=\mu_{\mathrm{bhv}}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, with a slight abuse of notation.
Additionally, the following standard (e.g., see [47, 48, 30]) assumption is adopted in our work. It can be ensured by normalizing the features.
Assumption 3.2
In this study, our primary focus is the estimation of state-action value functions, which we achieve by minimizing the Mean Squared Projected Bellman Error (MSPBE) through the application of smooth truncation operator:
$J(\theta)=\frac{1}{2}\mathbb{E}_{\mathcal{D}}\left(\Pi\mathcal{B}_{\tau,\delta}\hat{Q}_{\theta}(s,a)-\hat{Q}_{\theta}(s,a)\right)^{2}.$
Calculating the projected Bellman backup $\Pi\mathcal{B}_{\tau,\delta}\hat{Q}_{\theta}(s,a)$ presents significant challenges. This has motivated the development of the bi-level optimization formulation detailed below:
$\quad\min_{\theta\in\mathbb{R}^{d}}J\left(\theta\right)$ $\triangleq f\left(\theta,\omega^{*}(\theta)\right)$ (3)
$\triangleq\frac{1}{2}\mathbb{E}_{\mathcal{D}}\left(\phi(s,a)^{\top}\omega^{*}\left(\theta\right)-\phi(s,a)^{\top}\theta\right)^{2}$
@@ 3.2 Bi-level Formulation
In this context, $\omega$ represents the parameters of the main network, and $\theta$ represents the parameters of the target network.
In prior successful algorithms [49, 24, 50], the target network serves as a delayed replica of the main network, stabilizing the learning process and reducing overestimation. To achieve this, the target network’s parameters are updated less frequently or using smaller step sizes compared to the main network, enabling the target network to offer a slower-moving target for state-action value function estimates. The goal of updating the target deviates from simply mimicking the main network. Instead, the target network is updated with the objective of minimizing the MSPBE with smooth truncation. A key distinction in our application of the smooth truncation operator, compared to the approach in [32], lies in its implementation. We apply it to the state value function, whereas [32] employs it on the state-action value function. As a result, the policies corresponding to $Q\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ remain unchanged in the regularized MDP. We define the regularized policy $\pi_{\theta}$ corresponding to the approximated regularized state-action value function $\hat{Q}_{\theta}$ as
$\pi_{\theta}(\cdot\mid s)\triangleq\nabla G^{*}_{\tau}\left(\hat{Q}_{\theta}\left(s,\cdot\right)\right),\quad\forall s\in\mathcal{S}.$ (4)
The strong convexity of $g(\theta,\omega)$ in $\omega$ is established in the following proposition, which can be proved by using Taylor’s expansion.
Proposition 3.3
@@
The lower-level problem’s objective function aims to identify $\omega^{*}(\theta)$, an optimal weight of the approximation of the Bellman backup for a target network parameterized by $\theta$, within the subspace spanned by the basis functions. It is given by
@@
Under Assumption 3.1 and 3.2, the $\ell_{2}$-norm of $\omega^{*}(\theta)$ is bounded by $(R_{\max}+\gamma\delta)/\lambda_{g}$. In this case, the optimal value, $g(\theta,\omega^{*}(\theta))$, from the lower-level problem characterizes the square of the approximation error with respect to $\hat{Q}_{\theta}$ under the distribution $\mathcal{D}$. This error represents the component of $\mathcal{B}_{\tau,\delta}\hat{Q}_{\theta}$ that is not captured within the span of the basis functions. $\hat{Q}_{\omega^{*}(\theta)}=\Phi\omega^{*}(\theta)$ represents the projection of $\mathcal{B}_{\tau,\delta}\hat{Q}_{\theta}$ onto the space spanned by the basis functions with respect to the norm $\|\cdot\|_{D_{\mu_{\mathrm{bhv}}}}$, i.e. $\hat{Q}_{\omega^{*}(\theta)}=\Pi_{D_{\mu_{\mathrm{bhv}}}}\mathcal{B}_{\tau,\delta}\hat{Q}_{\theta}$. To simplify the presentation, we adopt a shorthand notation by using $\Pi$ to represent $\Pi_{D_{\mu_{\mathrm{bhv}}}}$ throughout the rest of this paper, albeit with a slight abuse of notation.
The optimization problem (3) can be solved by a gradient-based approach. The gradient of $J(\theta)$ can be derived by using the chain rule as
$\nabla J(\theta)$
$=$ $\nabla_{\theta}f(\theta,\omega^{*}(\theta))$
$-\nabla^{2}_{\theta\omega}g(\theta,\omega^{*}(\theta))\left[\nabla^{2}_{\omega\omega}g(\theta,\omega^{*}(\theta))\right]^{-1}\nabla_{\omega}f(\theta,\omega^{*}(\theta))$
@@
$\hat{\Sigma}_{\tau,\delta,\theta}\triangleq$
$\mathbb{E}_{\mathcal{D}}\!\!\left[\!\left(\!\!\gamma z\left(s^{\prime};\tau,\delta,\theta\right)\!\sum_{a^{\prime}\in\mathcal{A}}\!\!\!\pi_{\theta}\left(a^{\prime}\mid s^{\prime}\right)\phi\left(s^{\prime},a^{\prime}\right)-\phi(s,a)\!\!\right)\!\phi(s,a)^{\top}\!\right],$ (7)
@@
Observe that the range of $z$ is $(0,1)$. The following assumption is important in our analysis.
Assumption 3.4
@@
As the threshold $\delta$ approaches infinity, the smooth truncation operator $\mathcal{K}_{\delta}$ becomes equivalent to the identity function. Under these conditions, it is clear that $z\left(s;\tau,\delta,\theta\right)$ approaches $1$ for every $s\in\mathcal{S}$ given a finite value of $\theta$. Consequently, Assumption 3.4 is as strong as similar assumptions in [27, 51, 37, 23, 18, 52, 31]. Nonetheless, when $\delta$ has a finite value, Assumption 3.4 is weaker, especially for a small value of $\delta$.
It is computationally expensive to calculate $\nabla J(\theta)$ during implementation. Alternatively, we can employ a two-timescale algorithm to simultaneously update the parameters $\theta$ and $\omega$, each operating on a different timescale. For the upper-level problem, the surrogate gradient ¯$\bar{\nabla}_{\theta}f(\theta,\omega)$ is used with $\omega^{*}(\theta)$ replaced by $\omega$ in (3.2), defining the surrogate gradient as
¯$\bar{\nabla}_{\theta}f(\theta,\omega)=\hat{\Sigma}_{\tau,\delta,\theta}\left(\omega-\theta\right).$ (9)
In addition, the surrogate gradient can be readily sampled under a stochastic setting, similar to the gradient of the lower-level objective function, which is given by
$\nabla_{\omega}g(\theta,\omega)=\\ \mathbb{E}_{\mathcal{D}}\Big{[}\phi(s,a)\Big{(}\phi(s,a)^{\top}\omega-R(s,a)-\gamma\hat{G}_{\tau,\delta}(\hat{Q}_{\theta}\left(s^{\prime},\cdot\right))\Big{)}\Big{]}.$
One of the approaches for solving bi-level optimization problems is two-timescale algorithm. It simultaneously updates both the lower-level and upper-level solutions, with convergence guaranteed by employing a faster timescale for the lower-level solution and a slower timescale for the upper-level solution [53, 54, 55]. Within this framework, the target network, $\hat{Q}_{\theta}$, is updated at a slower pace compared to the main network, $\hat{Q}_{\omega}$, in our problem.
4 The Proposed Algorithm
@@
In this section, we introduce a gradient-based single-loop algorithm that extends the two-timescale stochastic approximation (TTSA) approach, as outlined in [55]. Our focus is on estimating the value functions of the regularized MDP with linear function approximation. We explore this problem within a framework characterized by Markovian noise, noting that our results can also be extended to settings with i.i.d. noise. We use $(s_{t},a_{t},s_{t}^{\prime})$ to represent the $t$-th transition on the Markov chain induced by $\pi_{\mathrm{bhv}}$ on MDP $\mathcal{M}_{\tau}$. The action $a_{t}$ is sampled from $\pi_{\mathrm{bhv}}(\cdot\mid s_{t})$ and the next state $s_{t+1}=s_{t}^{\prime}$, which is drawn from $P(\cdot\mid s_{t},a_{t})$. The stochastic gradient of the lower-level problem and the stochastic surrogate gradient of the upper-level problem are represented by
$h_{g}^{t}\triangleq\phi(s_{t},a_{t})\left(\phi(s_{t},a_{t})^{\top}\omega^{t}-R(s_{t},a_{t})-\gamma\hat{G}_{\tau,\delta}(\hat{Q}_{\theta^{t}}\left(s^{\prime}_{t},\cdot\right))\right)$ (10a)
@@
$h_{f}^{t}\triangleq\bigg{(}\gamma z\left(s^{\prime};\tau,\delta,\theta\right)\sum_{a^{\prime}\in\mathcal{A}}$ $\pi_{\theta^{t}}\left(a^{\prime}\mid s^{\prime}_{t}\right)\phi(s_{t}^{\prime},a^{\prime})-\phi(s_{t},a_{t})\bigg{)}$
×$\times\phi(s_{t},a_{t})^{\top}\left(\omega^{t+1}-\theta^{t}\right),$ (10b)
respectively. Recall the definition of $\pi_{\theta}$ in (4), which represents the optimal policy derived from the state-action value estimation of the target network, $\hat{Q}_{\theta}$. The parameters are updated with the step sizes $\alpha$ and $\beta$ in the following sense:
@@
where $\mathcal{P}$ denotes the projection onto $\ell_{2}$-balls with a radius of $(R_{\max}+\gamma\delta)/\lambda_{g}$. This value ensures that the optimal point $\omega^{*}(\theta^{t})$ of the lower-level problem lies within the specified ball. Utilizing the projection $\mathcal{P}$ constrains the main network, preventing it from taking overly large steps in the incorrect direction [48, 56, 52].
The update rule for the upper-level problem in Algorithm 1 sets it apart from existing two-timescale algorithms for off-policy value function estimation [27, 25, 28, 52, 31]. Specifically, it normalizes the stochastic gradient by dividing it by the difference between the weights of the main and target networks. Further analysis is provided in Section 5. See the two-timescale regularized Q-learning with adaptive stepsize in Algorithm 1.

### v3

@@ 2.3 Linear Function Approximation
Function approximation techniques are prevalent as they offer an effective solution to this problem by allowing agents to approximate the values using significantly fewer parameters. In this study, we focus on leveraging linear function approximation which approximates the value functions linearly in the given basis vectors.
The basis vectors $\phi_{i}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$, $i=1,2,\dots,d$ are, without loss of generality, linearly independent. We denote $\phi(s,a)=\left[\phi_{1}(s,a),\phi_{2}(s,a),\dots,\phi_{d}(s,a)\right]^{\top}\in\mathbb{R}^{d}$ for all $(s,a)$ and let $\Phi\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times d}$ defined by $\Phi=\left[\phi(s_{1},a_{1}),\dots,\phi(s_{|\mathcal{S}|},a_{|\mathcal{A}|})\right]^{\top}$. We approximate state-action function as $\hat{Q}_{\theta}\triangleq\Phi\theta$, where $\theta\in\mathbb{R}^{d}$ is the parameter vector. We also use $\hat{Q}_{\theta}(s,\cdot)=\phi(s,\cdot)^{\top}\theta$ to denote the vector of the approximated state-action values of all actions given the state $s\in\mathcal{S}$, where $\phi(s,\cdot)=\left[\phi(s,a_{1}),\dots,\phi(s,a_{|\mathcal{A}|})\right]$. By Proposition 2.2 (iv), the optimal policy satisfies:
$\pi_{\theta}(\cdot|s)$ $=\nabla G^{*}_{\tau}\big{(}\hat{Q}_{\theta}(s,\cdot)\big{)}$ (3)
Suppose $\mu$ is a distribution that exhibits full support across the state-action space, let $D_{\mu}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}||\mathcal{A}|}$ be a diagonal matrix with the distribution $\mu$ on its diagonal. Furthermore, we define the weighted $\ell_{2}$-norm with respect to the distribution $\mu$ as $\|x\|_{D_{\mu}}\triangleq\sqrt{x^{\top}D_{\mu}x}$. Let $\Pi_{D_{\mu}}$ be the projection operator which maps a vector onto the span of the basis functions with respect to the norm $\|\cdot\|_{D_{\mu}}$; it is given by $\Pi_{D_{\mu}}=\Phi\left(\Phi^{\top}D_{\mu}\Phi\right)^{-1}\Phi^{\top}D_{\mu}$ under the assumption that $\Phi^{\top}D_{\mu}\Phi$ is a positive definite matrix.
@@
2.4 Standing Assumptions
We now state the standing assumptions for our analysis. Let $\pi_{\mathrm{bhv}}$ be the behavioral policy used to collect data in regularized MDP $\mathcal{M}_{\tau}$ and we adopt the following assumption.
Assumption 2.2.
Assumption 2.2 guarantees the ergodicity of the Markov chain, which in turn ensures that it converges to its stationary distribution at a geometric rate. Furthermore, Let $\mathcal{D}$ be the stationary distribution of the transition tuples induced by $\pi_{\mathrm{bhv}}$. Consequently, $\mathcal{D}=\mu_{\mathrm{bhv}}\otimes P$, where $\otimes$ denotes the tensor product between two distributions. We use the term $\mathcal{D}$ to also refer to the marginal distribution of state-action pairs, i.e., $\mathcal{D}(s,a)=\mu_{\mathrm{bhv}}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, with a slight abuse of notation. Finally we make the following standing assumption
Assumption 2.3.
[…]
@@ 3.1 Smooth Truncation Operator
In addition, we let $\mathcal{B}_{\tau,\delta}$ be the smooth truncated optimal regularized Bellman operator which is given as, for an arbitrary mapping $Q:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$:
$\mathcal{B}_{\tau,\delta}Q(s,a)\triangleq R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot\mid s,a)}\left[\hat{G}_{\tau,\delta}\left(Q\left(s^{\prime},\cdot\right)\right)\right],$
@@
3.2 A Bi-level Formulation for Bellman Error Minimization
In what follows, we aim to minimize the Mean Squared Projected Bellman Error (MSPBE) defined as
$J(\theta)\triangleq\frac{1}{2}\mathbb{E}_{\mathcal{D}}\left(\Pi\mathcal{B}_{\tau,\delta}\hat{Q}_{\theta}(s,a)-\hat{Q}_{\theta}(s,a)\right)^{2}.$
A bi-level optimization formulation for minimizing MSPBE is as follows:
$\quad\min_{\theta\in\mathbb{R}^{d}}f\left(\theta,\omega^{*}(\theta)\right)$ $\triangleq\frac{1}{2}\mathbb{E}_{\mathcal{D}}\left(\phi(s,a)^{\top}\omega^{*}\left(\theta\right)-\phi(s,a)^{\top}\theta\right)^{2}$ (5)
@@ 3.2 A Bi-level Formulation for Bellman Error Minimization
Note that $\hat{Q}_{\omega^{*}(\theta)}=\Phi\omega^{*}(\theta)$ represents the projection of $\mathcal{B}_{\tau,\delta}\hat{Q}_{\theta}$ onto the space spanned by the basis functions with respect to the norm $\|\cdot\|_{D_{\mu_{\mathrm{bhv}}}}$, i.e. $\hat{Q}_{\omega^{*}(\theta)}=\Pi\mathcal{B}_{\tau,\delta}\hat{Q}_{\theta}$. Hence, it can be readily seen that the upper level objective function is MSPBE, i.e. $J\left(\theta\right)=f\left(\theta,\omega^{*}(\theta)\right)$. In the above formulation, the parameters $\omega$ are referred to as the main solution (or network) whereas $\theta$ represents the parameters of the target solution (or network).
We now show the lower level objective function $g(\theta,\omega)$ is strongly convex in $\omega$:
Proposition 3.1.
@@
See the proof in Appendix E.1. In addition, the solution $\omega^{*}(\theta)$ of the lower level problem has the following closed-form:
@@
3.3 Gradient of Bellman Error
The gradient of $J(\theta)$ can be derived by using the chain rule as
$\nabla J(\theta)$ $=\nabla_{\theta}f(\theta,\omega^{*}(\theta))-\nabla^{2}_{\theta\omega}g(\theta,\omega^{*}(\theta))\left[\nabla^{2}_{\omega\omega}g(\theta,\omega^{*}(\theta))\right]^{-1}\nabla_{\omega}f(\theta,\omega^{*}(\theta))$
@@ 3.3 Gradient of Bellman Error
$\hat{\Sigma}_{\tau,\delta,\theta}\triangleq\mathbb{E}_{\mathcal{D}}\Big{[}\Big{(}\gamma z\left(s^{\prime};\tau,\delta,\theta\right)\sum_{a^{\prime}\in\mathcal{A}}\pi_{\theta}\left(a^{\prime}\mid s^{\prime}\right)\phi\left(s^{\prime},a^{\prime}\right)-\phi(s,a)\Big{)}\phi(s,a)^{\top}\Big{]},$
@@
Observe that the range of $z$ is $(0,1]$. We make the following regularity assumption on the gradient:
Assumption 3.2.
@@
Note that Assumption 3.2 is as strong as similar assumptions in [20, 36, 22, 37, 38, 39, 33]. This is because $z\left(s;\tau,\delta,\theta\right)\rightarrow 1$ as $\delta\rightarrow+\infty$. Nonetheless, when $\delta$ has a finite value, Assumption 3.2 is weaker, especially for a small value of $\delta$.
4 A Single-Loop Algorithm
In this section, we introduce a single-loop algorithm for solving the bi-level problem (5) wherein at every iteration the upper level and the lower solutions are updated based upon samples from a surrogate gradient of the upper level objective and the gradient of the lower level objective. Specifically, the gradient of the lower level objective is:
$\nabla_{\omega}g(\theta,\omega)$ $=\mathbb{E}_{\mathcal{D}}\big{[}\phi(s,a)\big{(}\phi(s,a)^{\top}\omega-R(s,a)-\gamma\hat{G}_{\tau,\delta}(\hat{Q}_{\theta}\left(s^{\prime},\cdot\right))\big{)}\big{]}$
A surrogate gradient for the upper level objective:
¯$\bar{\nabla}_{\theta}f(\theta,\omega)\triangleq\hat{\Sigma}_{\tau,\delta,\theta}\left(\omega-\theta\right).$ (9)
wherein $\omega^{*}(\theta)$ is replaced by $\omega$ in (3.3). Samples for the upper-level surrogate gradient and lower-level gradient can be readily obtained as follows.
@@ 4 A Single-Loop Algorithm
Let $(s_{t},a_{t},s_{t}^{\prime})$ denote the $t$-th transition on the Markov chain induced by $\pi_{\mathrm{bhv}}$ on MDP $\mathcal{M}_{\tau}$. The action $a_{t}$ is sampled from $\pi_{\mathrm{bhv}}(\cdot\mid s_{t})$ and the next state $s_{t+1}=s_{t}^{\prime}$, which is drawn from $P(\cdot\mid s_{t},a_{t})$. The stochastic gradient of the lower-level problem and the stochastic surrogate gradient of the upper-level problem are represented by:
$h_{g}^{t}\triangleq\phi(s_{t},a_{t})\left(\phi(s_{t},a_{t})^{\top}\omega^{t}-R(s_{t},a_{t})-\gamma\hat{G}_{\tau,\delta}(\hat{Q}_{\theta^{t}}\left(s^{\prime}_{t},\cdot\right))\right)$ (10)
@@
$h_{f}^{t}\triangleq\Big{(}\gamma z_{t}\sum_{a^{\prime}\in\mathcal{A}}$ $\pi_{\theta^{t}}\left(a^{\prime}\mid s^{\prime}_{t}\right)\phi(s_{t}^{\prime},a^{\prime})-\phi(s_{t},a_{t})\Big{)}$
×$\times\phi(s_{t},a_{t})^{\top}\left(\omega^{t+1}-\theta^{t}\right),$ (11)
respectively, where $z_{t}:=z\big{(}s_{t}^{\prime};\tau,\delta,\theta^{t}\big{)}$. Recall the definition of $\pi_{\theta}$ in (3), which represents the optimal policy derived from the state-action value estimation of the target network, $\hat{Q}_{\theta}$. The parameters are updated with the step sizes $\alpha$ and $\beta$ in the following sense:
@@
where $\mathcal{P}$ denotes the projection onto $\ell_{2}$-balls with a radius of $(R_{\max}+\gamma\delta)/\lambda_{g}$. This value ensures that the optimal point $\omega^{*}(\theta^{t})$ of the lower-level problem lies within the specified ball. The algorithm for regularized Q-learning with adaptive stepsize is summarized in Algorithm 1.
Remark: Utilizing the projection $\mathcal{P}$ constrains the main network, preventing it from taking overly large steps in the incorrect direction [35, 40, 39]. The update rule for the upper-level problem in Algorithm 1 normalizes the stochastic gradient by dividing it by the difference between the weights of the main and target networks.
Lemma 4.1.
(Asymptotic stability.) Suppose Assumptions 2.2, 2.3 and 3.2 hold. The approximated Q-value remains bounded almost surely for all $a\in\mathcal{A}$:
[…]
See the proof in Appendix C.1.